6. Suppose the demand and supply curves for a particular product are given below.

D: P = 160 − Q

S: P = 10 + 2Q

What is the equilibrium price, Pe , and equilibrium, Qe ?

a. Pe = 140; Qe = 10.

b. Pe = 80; Qe = 50.

c. Pe = 110; Qe = 50.

d. Pe = 50; Qe = 40.

e. None of the above.

Please show detail.
Answer #1

Option c.

The equilibrium price and quantity is found where demand and supply gets equated

D=S

160-Q=10+2Q

150=3Q

Q=50

Now by substituting Q=50 in demand equation, we get, D=160-Q=160-50=110